George is 5 years older than Mary. The sum of their ages is 43 years. How old are George and Mary?
Answer by algebrahouse.com(1659)

x = Mary's age
x + 5 = George's age

x + x + 5 = 43 {sum of ages is 43}
2x + 5 = 43 {combined like terms}
2x = 38 {subtracted 5 from each side}
x = 19 {divided each side by 2}
x + 5 = 24 {substituted 19, in for x, into x + 5}

Mary is 19
George is 24

For more help from me, visit: www.algebrahouse.com